OpenStudy (bloomlocke367):

Suppose you invest $500 in a savings account that pays 3.5% annual interest. When will the account contain at least $650?

OpenStudy (bloomlocke367):

@amistre64

OpenStudy (bloomlocke367):

I have the equation \(A(t)=a(1+2)^t\) that I believe I need to use.

OpenStudy (bloomlocke367):

oops, that 2 should be r

OpenStudy (bloomlocke367):

so I think a is 500 and r would be 0.035... and I think I'm solving for t...

Nnesha (nnesha):

what is a in the question what is r ?

Nnesha (nnesha):

:-) yep right

OpenStudy (bloomlocke367):

a is the initial amount and r is the rate of growth, or the rate of decay

OpenStudy (bloomlocke367):

and A would be 650?

Nnesha (nnesha):

okay there is two ways you can solve this problem first one you can replace t by 2 3 or 5 6 and see which one give at least 650 when you solve right side another way replace a(t) by 650

OpenStudy (bloomlocke367):

so do I need to make a table?

Nnesha (nnesha):

nope for example first method i'll replace a by 500 and r by .035 and try 2 for t \[a(t)= 500(1+.035)^2\] like this now i will solve right side if i get at least 650 then 2 is my answer

OpenStudy (bloomlocke367):

so it's a guess and check thing? would a table work?

OpenStudy (bloomlocke367):

and 2 doesn't work.

Nnesha (nnesha):

yep but i guess you should do by 2nd method :) \[\huge\rm 650 = 500(1+.035)^t\]

Nnesha (nnesha):

yep guessing for first one :)

Nnesha (nnesha):

yep but i guess you should do by 2nd method :) \[\huge\rm 650 = 500(1+.035)^t\] do you know how to solve this one ?

OpenStudy (bloomlocke367):

is there a way to solve it algebraically to find t?

Nnesha (nnesha):

do you know how to take ln ?? log ?

OpenStudy (bloomlocke367):

No, that's next. This came first...

Nnesha (nnesha):

okay then guess :)

OpenStudy (bloomlocke367):

The only way I know how to do this is to make a table... :/

OpenStudy (bloomlocke367):

okay.

Nnesha (nnesha):

yep that one :-)

OpenStudy (bloomlocke367):

I get 8.. because it equals 658.4045...

OpenStudy (bloomlocke367):

so after 8 years.

Nnesha (nnesha):

:-)

OpenStudy (bloomlocke367):

Can you help with some more?

Nnesha (nnesha):

i'll try :)